OpenStudy (anonymous):

Solve -8d- 9= -3(d+3) -5d

OpenStudy (anonymous):

use foil for the brackets -8d-9=-3d-9-5d

OpenStudy (anonymous):

-8d-9=-3(d+3)-5d -8d-9=-3d-9-5d ( arrange term and solve it )